If 0.1 mole of COC1₂, in a 1- liter vessel is 60% dissociated into CO and Cl, at a certain temperature, calculate the equilibrium constant, Kc.


1
Expert's answer
2021-10-30T01:15:10-0400

Initially, 2 moles of CO are present.


At equilibrium, 1 mole of CO is present.


Hence, 2−1=1 mole of CO has reacted.


1 mole of CO will react with 1 mole of Cl 2 to form 1 mole of COCl2

​3−1=2 moles of Cl2​ remains at equilibrium.


The equilibrium constant

"K_c= \frac{\frac{1mol}{5L}}{\frac{1mol}{5L}×\frac{2mol}{5L}}"


"K_c= 2.5"